# GeoLangBind: Unifying Earth Observation with Agglomerative Vision–Language Foundation Models

### v1

@@ Abstract
Earth observation (EO) data, collected from diverse sensors with varying imaging principles, present significant challenges in creating unified analytical frameworks. We present GeoLangBind, a novel agglomerative vision–language foundation model that bridges the gap between heterogeneous EO data modalities using language as a unifying medium. Our approach aligns different EO data types into a shared language embedding space, enabling seamless integration and complementary feature learning from diverse sensor data. To achieve this, we construct a large-scale multimodal image–text dataset, GeoLangBind-2M, encompassing six data modalities. GeoLangBind leverages this dataset to develop a zero-shot foundation model capable of processing arbitrary numbers of EO data channels as input. Through our designed Modality-aware Knowledge Agglomeration (MaKA) module and progressive multimodal weight merging strategy, we create a powerful agglomerative foundation model that excels in both zero-shot vision–language comprehension and fine-grained visual understanding. Extensive evaluation across 23 datasets covering multiple tasks demonstrates GeoLangBind’s superior performance and versatility in EO applications, offering a robust framework for various environmental monitoring and analysis tasks. The dataset and pretrained models will be publicly available.
1 Introduction
Figure 1: GeoLangBind leverages language as a bridge to unify heterogeneous remote sensing data, enabling a common representation space across diverse modalities for improved multimodal learning.
@@ 1 Introduction
Earth observation (EO) technologies have undergone remarkable advancement in recent years, generating vast amounts of multimodal geospatial data spanning different sensors and scales [54, 55, 70, 59]. In response to the exponential growth of remote sensing data, researchers have proposed foundation models pretrained on large-scale EO data, including SatMAE [12], Scale-MAE [38], HyperSIGMA [46], SSL4EO [47, 39], and SpectralGPT [20]. However, these models typically focus on a single EO data modality, limiting their ability to leverage complementary characteristics across different sensor types.
@@
To address this limitation, recent works such as OFA-Net [58], DOFA [57], AnySat [5], OmniSat [6], and FoMo-Net [8] have introduced foundation models capable of learning from multiple EO modalities within a unified framework. These approaches have demonstrated promising results in multimodal EO representation learning. However, their focus remains on unsupervised visual pretraining, lacking a structured way to align EO data with language for better interpretability and retrieval. Meanwhile, multimodal large language models like GPT [33], LLaMA [44], LLaVA [29], and DeepSeek [27] have shown remarkable capabilities in handling diverse and complex tasks. Vision-language foundation models, such as CLIP [36] and SigLIP [64], have played a crucial role as vision encoders of such multimodal large language models. Inspired by these advances, EO-specific adaptations like RemoteCLIP [28], RS5M [67], and SkyScript [50] have attempted to bring contrastive learning to remote sensing. However, these models are primarily limited to high-resolution RGB optical imagery, and can not handle other diverse EO modalities (e.g., multispectral data, Synthetic Aperture Radar data) that are critical for many applications.
Training a single vision–language foundation model capable of supporting multiple heterogeneous EO data modalities presents several challenges: 1) the lack of comprehensive multimodal EO datasets with aligned textual descriptions, 2) the design of a flexible architecture to handle variable-length input channels across different modalities, and 3) addressing the imbalance among different modalities in the training data. To address such challenges, we propose GeoLangBind, a unified multimodal vision–language foundation model that handles diverse data modalities and aligns them to the shared language feature space, as illustrated in Fig. 1. In addition, GeoLangBind enhances fine-grained semantic understanding via knowledge agglomeration [37] from multiple teacher models beyond contrastive learning.
To achieve this, we combine the following key designs: 1) a wavelength-aware dynamic encoder to handle multiple data modalities, 2) a modality-aware knowledge agglomeration module, and 3) progressive weight-space merging [53] to scale the model efficiently across multiple modalities. Our contributions are as follows:
@@
    We construct GeoLangBind-2M, a comprehensive dataset of two million image–text pairs spanning six distinct EO data modalities to facilitate the training of multimodal geospatial foundation models.
@@
    We propose an effective Modality-aware Knowledge Agglomeration (MaKA) module to enhance continuous pre-training and improve fine-grained image understanding.
@@
    We design a progressive weight merging strategy that enables scalable model adaptation across multiple modalities, outperforming traditional data mixing approaches.
@@
    We conduct extensive experiments on zero-shot classification, semantic segmentation, and cross-modal image retrieval, demonstrating the superiority of GeoLangBind across diverse tasks.
Benchmarking results indicate that GeoLangBind achieves state-of-the-art performance while being flexible in handling diverse data modalities with any number of spectral bands.
Figure 2: Visualization of data samples from our GeoLangBind-2M. The dataset includes imagery from six different sensors and modalities: Sentinel-2 multispectral, Sentinel-1 SAR, EnMAP hyperspectral, elevation maps, infrared imagery, and aerial imagery. Each sample is paired with textual descriptions capturing key land cover types, objects, and geographic features.
2 Related work
@@
Existing pretrained models predominantly focus on RGB data, such as GFM [31], Scale-MAE [38], and Cross-Scale MAE [43]. Others specialize in multispectral imagery, with FG-MAE [48] and SatMAE [12] tailored for Sentinel-2 data, while SSL4EO-L [39] is trained on Landsat imagery. SpectralGPT [20] leverages a 3D generative transformer for spectral data analysis. Beyond optical data, CROMA [14] introduces dual unimodal encoders for multispectral and Synthetic Aperture Radar (SAR) data, using a cross-modal radar–optical transformer to learn unified deep representations. DeCUR [49] enhances bimodal learning by decoupling unique and shared features across modalities. Meanwhile, Satlas [7] assembles a large-scale multi-sensor dataset, pretraining separate models for each sensor. DOFA [57] introduces a neural plasticity-inspired hypernetwork that adapts to different sensor wavelengths, enabling joint training on five distinct sensors. OmniSat [6] proposes to fuse features across multiple EO modalities to learn robust multimodal representations without labeled data. SkySense [18] utilizes separate visual encoders for different data modalities and designs a dedicated module for multimodal fusion.
Zero-shot foundation models
Zero-shot foundation models like CLIP [36] have significantly advanced computer vision by enabling models to generalize across tasks without explicit task-specific training. Building upon CLIP, SigLIP [64] introduces a pairwise sigmoid loss function to enlarge the training batch and enhance the model performance. Zero-shot foundation models have also been investigated in EO applications. SkyScript [50] constructs a large-scale image–text dataset by linking remote sensing images with OpenStreetMap [34] semantics. RemoteCLIP [28] introduces an image–text dataset by generating captions from existing annotated datasets. GeoRSCLIP [67] introduces RS5M, an image–text dataset with 5M pairs, created by filtering existing datasets and generating captions for label-only remote sensing data. Despite these advances, existing vision–language models mainly focus on RGB data and have limitations in learning representations from multiple EO data modalities, which are crucial for holistic environmental analysis. To unify multiple data modalities, ImageBind [16] aligns six modalities within a shared embedding space. LanguageBind [69] extends this by using language as a universal intermediary for modality alignment. While these approaches demonstrate the potential of language or vision as unifying mediums, applying them to EO data is still challenging due to the scarcity of multimodal image–text datasets and the diverse scales of Earth systems.
3 Dataset construction: GeoLangBind-2M
Table 1: Summary of dataset sources of GeoLangBind-2M. * indicates image–text datasets that are created in this work. SAR = synthetic aperture radar, MSI = multispectral imagery, HSI = hyperspectral imagery, and IR = infrared.
@@ 3 Dataset construction: GeoLangBind-2M
GeoLangBind-2M contains a broad range of data modalities including RGB, SAR, multispectral, hyperspectral, infrared, and elevation data, as listed in Tab. 1. Each image is paired with a high-quality text annotation, as illustrated in Fig. 2. All datasets are implemented using the TorchGeo library [40]. In the following, we introduce the composition of the dataset.
@@
High-resolution RGB images with textual descriptions form a significant portion of GeoLangBind-2M. We incorporate parts of the RemoteCLIP [28] dataset, i.e. Seg4 and Det10, which are constructed by generating captions from existing segmentation and detection datasets. We also include VRSBench [26], which aggregates multiple object detection datasets. SkyScript [50] is a large-scale image–text dataset for EO, from which we include the top 30% highest quality samples after filtering. Beyond existing datasets, we introduce Flair2-RGB-caption [15], a newly created image–text dataset. Using semantic information as prompts, we employ the Pixtral-12B [2] model to generate captions for detailed land cover features.
@@
To enhance the dataset with SAR–text pairs, we create MMFlood-SAR-caption [32], which contains captions emphasizing water extent and flooded regions. Additionally, the SAR-ship-caption dataset includes captions describing ships in SAR images. ChatEarthNet-SAR-caption further extends the collection by covering diverse SAR scenes from Sentinel-1, focusing on land cover types.
@@
We construct ChatEarthNet-S2-caption by summarizing long, detailed descriptions from the ChatEarthNet dataset [63]. This dataset contains Sentinel-2 multispectral imagery covering more than 10 land cover types. Furthermore, we introduce NLCD-hyper-caption, a hyperspectral dataset with 200+ EnMAP bands [17]. Like Flair2-RGB-caption, we use Pixtral-12B to generate rich captions, leveraging annotated land cover maps for enhanced descriptiveness. Similarly, we also generate captions using Pixtral-12B to construct the Flair2-Elevation-caption dataset. Further details on the construction of GeoLangBind-2M are provided in the Supplementary materials.
@@
Figure 3: GeoLangBind aims to unify and align diverse Earth observation data modalities into the shared language feature space while enhancing fine-grained image understanding. The model uses modality-aware knowledge agglomeration (MaKA) to transfer knowledge from multiple teachers. The wavelength-aware dynamic patch embedding module processes multimodal images dynamically.
4 Methodology
4.1 GeoLangBind model design
GeoLangBind aims to unify diverse data modalities and embed them into a shared language feature space while enhancing fine-grained image understanding ability. To this end, we design a wavelength-aware dynamic encoder, a modality-aware knowledge agglomeration module, and a progressive weight-space merging approach.
4.2 Wavelength-aware dynamic encoder
@@
To handle the diversity of spectral bands across modalities, we adopt the dynamic encoder architecture from DOFA [57], as illustrated in Fig. 3. Given an input image $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$ with $C$ spectral channels, we define its corresponding central wavelengths as $\mathbf{\lambda}\in\mathbb{R}^{C}$. For modalities without well-defined wavelengths, such as elevation, we assign default wavelengths equivalent to those of the RGB channels. First, a 1D sine-cosine positional encoding is applied to embed these wavelengths into a higher-dimensional space: $\mathbf{V}_{\lambda}=\text{PE}(\mathbf{\lambda})\in\mathbb{R}^{C\times d_{\lambda}},$ where $d_{\lambda}$ is the embedding dimension. The encoded wavelengths $\mathbf{V}_{\lambda}$ are then transformed using two fully-connected layers with residual connections.
Next, a lightweight Transformer [45] encoder processes the transformed wavelength embeddings. Specifically, we concatenate the transformed embeddings $\mathbf{V^{\prime}}_{\lambda}$ with the learnable weight query tokens $\mathbf{Q}_{w}$ and a learnable bias query token $\mathbf{Q}_{b}$ as input to the Transformer encoder. After the Transformer layer, the output features are passed through two fully-connected layers to generate the dynamic convolutional weights $\mathbf{M}_{w}$ and biases $\mathbf{M}_{b}$. Finally, these dynamically generated weights and biases are used for patch embedding (implemented as convolution). This process enables the model to handle different data modalities with varying spectral channels using a unified deep encoder.
4.3 Modality-aware knowledge agglomeration
@@
We introduce the modality-aware knowledge agglomeration (MaKA) module, which leverages wavelength as a modality-specific condition when distilling features from different data modalities. MaKA comprises three key components: a wavelength-aware prompt generator, which encodes wavelength values into a learnable embedding vector; a modality-aware conditional layer normalization, which adapts normalization parameters based on modality-specific characteristics; and a feature projector, which transforms input features to facilitate effective feature matching and knowledge distillation. Below, we introduce each of them in detail.
Wavelength-aware prompt generator

### v2

# DOFA-CLIP: Multimodal Vision–Language Foundation Models for Earth Observation

@@ Abstract
Earth observation (EO) spans a broad spectrum of modalities, including optical, radar, multispectral, and hyperspectral data, each capturing distinct environmental signals. However, current vision-language models in EO, particularly CLIP-based variants, remain confined to individual modalities, limiting generalization and scalability across diverse tasks. We present DOFA-CLIP (Dynamic-One-For-All CLIP), a unified vision-language foundation model that dynamically adapts to EO modalities with flexible spectral configurations through a single Transformer backbone. Our approach introduces three key contributions: 1) the construction of GeoLangBind-2M, a large-scale EO image–text dataset covering six heterogeneous modalities with rich natural language descriptions; 2) a novel training strategy called VECT (Vision-models Enhanced Contrastive Text-image pretraining), which enhances the spatial awareness of CLIP features with multiple vision foundation models; and 3) a Modality-aware Knowledge Agglomeration (MaKA) module that refines feature distillation with modality-specific awareness. DOFA-CLIP achieves state-of-the-art zero-shot performance across a wide range of EO benchmarks, including unseen modalities and a diverse number of input spectral bands. Together, these contributions establish a scalable foundation for multimodal EO understanding and open new avenues for integrating heterogeneous EO data with large language models. Code and datasets will be released.
Introduction
Figure 1: DOFA-CLIP employs a single unified vision Transformer to align heterogeneous EO data modalities using language as the grounding signal for cross-modal understanding and zero-shot generalization.
@@ Introduction
Earth observation (EO) technologies have undergone remarkable advancement in recent years, generating vast amounts of multimodal geospatial data spanning different sensors and scales (Xia et al. 2018; Zhu et al. 2017; Xiong et al. 2024c). In response to the exponential growth of remote sensing data, researchers have proposed foundation models (Cong et al. 2022; Reed et al. 2023; Wang et al. 2024a, 2023a; Stewart et al. 2023; Hong et al. 2024) pretrained on large-scale EO data for multiple downstream tasks. However, these models typically focus on individual EO data modality, limiting their ability to leverage complementary characteristics across different sensor types (Zhu et al. 2024).
@@
To address this limitation, recent works such as OFA-Net (Xiong et al. 2024b), DOFA (Xiong et al. 2024a), AnySat (Astruc et al. 2024a), OmniSat (Astruc et al. 2024b), and FoMo-Net (Bountos, Ouaknine, and Rolnick 2023) have introduced foundation models capable of learning from multiple EO modalities within a unified framework. These approaches have demonstrated promising results in multimodal EO representation learning. However, their focus remains on unsupervised visual pretraining, lacking a structured way to align EO data with language for better interpretability and retrieval. Meanwhile, multimodal large language models like GPT (OpenAI 2023), LLaVA (Liu et al. 2023), and Pixtral (Agrawal et al. 2024) have shown remarkable capabilities in handling diverse and complex tasks. Vision-language foundation models, such as CLIP (Radford et al. 2021) and SigLIP (Zhai et al. 2023), have played a crucial role as vision encoders of such multimodal large language models. Inspired by these advances, EO-specific adaptations like RemoteCLIP (Liu et al. 2024), RS5M (Zhang et al. 2024b), RS-CLIP (Li et al. 2023), and SkyScript (Wang et al. 2024c) have attempted to bring vision-language contrastive learning to remote sensing. However, existing CLIP models for EO have limitations: 1) These models are usually limited to high-resolution RGB imagery and cannot handle modalities like multi/hyper-spectral or SAR, which are vital for many applications. 2) Compared with vision models like DINOv2 (Oquab et al. 2023), CLIP models are unsuitable for dense understanding tasks due to the lack of spatial awareness.
Training a single powerful vision–language model capable of handling heterogeneous EO data modalities raises several core research questions (RQ). These include:
[…]
    RQ1: How can we build a large-scale, semantically rich image–text dataset for diverse heterogeneous EO modalities?
[…]
    RQ2: How can we efficiently train large CLIP models for EO with better spatial awareness?
[…]
    RQ3: How can a single CLIP model handle heterogeneous EO data while preserving modality-specific characteristics?
To address such challenges, we propose Dynamic-One-For-All CLIP (DOFA-CLIP), a unified multimodal vision–language foundation model that aligns diverse EO modalities to the language space using a single Transformer, as shown in Fig. 1. Our contributions are as follows:
@@
    We introduce GeoLangBind-2M, a large-scale EO dataset comprising two million image–text pairs across six heterogeneous modalities, providing high-quality text descriptions for multimodal vision–language training.
@@
    We propose Vision-models Enhanced Contrastive Text-image pretraining (VECT), a simple yet effective training strategy that enhances the fine-grained spatial awareness of CLIP models with open vision foundation models.
@@
    We develop the Modality-aware Knowledge Agglomeration (MaKA) module to support modality-adaptive representation distillation within a unified Transformer.
@@
    We conduct extensive evaluations across zero-shot classification and cross-modal retrieval tasks, demonstrating state-of-the-art performance and better spatial awareness on both seen and unseen EO modalities.
Figure 2: Visualization of data samples from GeoLangBind-2M. The dataset includes imagery from six different sensors and modalities: Sentinel-2 multispectral, Sentinel-1 SAR, EnMAP hyperspectral, elevation maps, infrared imagery, and aerial imagery. Each sample is paired with textual descriptions capturing key land cover types, objects, and geographic features.
Related work
@@
Existing pretrained models predominantly focus on RGB data, such as GFM (Mendieta et al. 2023), Scale-MAE (Reed et al. 2023), and Cross-Scale MAE (Tang et al. 2024). Others specialize in multispectral imagery, with FG-MAE (Wang et al. 2023b) and SatMAE (Cong et al. 2022) tailored for Sentinel-2 data, while SSL4EO-L (Stewart et al. 2023) is trained on Landsat imagery. SpectralGPT (Hong et al. 2024) leverages a 3D generative transformer for spectral data analysis. CROMA (Fuller, Millard, and Green 2023) and DeCUR (Wang et al. 2024b) use self-supervised multimodal learning to learn features across modalities. Satlas (Bastani et al. 2023) assembles a large-scale multi-sensor dataset, pretraining separate models for each sensor. DOFA (Xiong et al. 2024a) introduces a neural plasticity-inspired hypernetwork that adapts to different sensor wavelengths, enabling joint training on five distinct sensors. OmniSat (Astruc et al. 2024b) proposes to fuse features across multiple EO modalities to learn robust multimodal representations without labeled data. SkySense (Guo et al. 2024) utilizes separate visual encoders for different data modalities and designs a dedicated module for multimodal fusion.
Vision-language foundation models
Zero-shot foundation models like CLIP (Radford et al. 2021) have significantly advanced computer vision. Building upon CLIP, SigLIP (Zhai et al. 2023) introduces a pairwise sigmoid loss function to enlarge the training batch and enhance the model performance. Zero-shot foundation models have also been investigated in EO applications. SkyScript (Wang et al. 2024c) constructs a large-scale image–text dataset by linking remote sensing images with OpenStreetMap (OpenStreetMap contributors 2017) semantics. RemoteCLIP (Liu et al. 2024) introduces an image–text dataset by generating captions from existing annotated datasets. GeoRSCLIP (Zhang et al. 2024b) introduces RS5M, an image–text dataset with 5M pairs, created by filtering existing datasets and generating captions for label-only remote sensing data. Despite these advances, existing vision–language models mainly focus on RGB data and have limitations in learning representations from multiple EO data modalities. To unify multiple data modalities, ImageBind (Girdhar et al. 2023) and TaxaBind (Sastry et al. 2025) align multiple modalities within a shared embedding space. LanguageBind (Zhu et al. 2023) uses language as a universal intermediary for modality alignment. In the era of Large Language Models (LLMs), methods such as GeoChat (Kuckreja et al. 2024), SkyEyeGPT (Zhan, Xiong, and Yuan 2025), and TeoChat (Irvin et al. 2024) integrate CLIP-based vision models with LLMs to unify diverse tasks and achieve impressive performance.
GeoLangBind-2M Dataset
Table 1: Summary of dataset sources of GeoLangBind-2M. * indicates image–text datasets created in this work. SAR = synthetic aperture radar, MSI = multispectral imagery, HSI = hyperspectral imagery, and IR = infrared.
@@
In response to RQ1, we propose GeoLangBind-2M, a large-scale image-text dataset covering a broad range of modalities including RGB, SAR, multispectral, hyperspectral, infrared, and elevation data, as listed in Table 1. Each image is paired with a high-quality text annotation, as illustrated in Fig. 2. All datasets are implemented using the TorchGeo library (Stewart et al. 2022). In the following, we introduce the composition of the dataset.
@@
High-resolution RGB images with textual descriptions form a significant portion of GeoLangBind-2M. We incorporate parts of the RemoteCLIP (Liu et al. 2024) dataset: Seg4 and Det10. We also include the training split of VRSBench (Li, Ding, and Elhoseiny 2024), which aggregates multiple object detection datasets. SkyScript (Wang et al. 2024c) is a global image–text dataset for EO, from which we include the top 30% highest quality samples. In addition to existing datasets, we construct the Flair2-RGB-caption (Garioud et al. 2023) dataset by leveraging the Pixtral-12B (Agrawal et al. 2024) model, using semantic information as prompts to generate captions for land cover features.
@@
To enhance the dataset with SAR–text pairs, we create MMFlood-SAR-caption (Montello, Arnaudo, and Rossi 2022), which contains captions emphasizing water extent and flooded regions (more details in supplementary material). Additionally, the SAR-ship-caption dataset includes captions describing ships in SAR images. ChatEarthNet-SAR-caption further extends the collection by covering diverse SAR scenes from Sentinel-1, focusing on land cover types.
@@
We construct ChatEarthNet-S2-caption by summarizing long, detailed descriptions from the ChatEarthNet dataset (Yuan et al. 2024). This dataset contains Sentinel-2 multispectral imagery covering more than 10 land cover types. Furthermore, we introduce NLCD-hyper-caption (Braham et al. 2024), a hyperspectral dataset with 200+ EnMAP bands (Guanter et al. 2015). Like Flair2-RGB-caption, we use Pixtral-12B to generate rich captions, leveraging annotated land cover maps for enhanced descriptiveness. Similarly, we also generate captions using Pixtral-12B to construct the Flair2-Elevation-caption dataset. Further details on the construction of GeoLangBind-2M are provided in the Supplementary materials.
@@
Figure 3: DOFA-CLIP aims to align diverse Earth observation data modalities into the language feature space. Vision-models Enhanced Contrastive Text-image pretraining (VECT) is proposed to enhance fine-grained image understanding. Modality-aware knowledge agglomeration (MaKA) is designed to transfer knowledge from heterogeneous sensors. The wavelength-aware dynamic patch embedding module processes multimodal images dynamically.
Dynamic-One-For-All CLIP
Wavelength-aware dynamic encoder
@@
To handle the diversity of spectral bands across modalities, we adopt the dynamic encoder architecture from DOFA (Xiong et al. 2024a), as illustrated in Fig. 3. Given an input image $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$ with $C$ spectral channels, we define its corresponding central wavelengths as $\mathbf{\lambda}\in\mathbb{R}^{C}$. For modalities without well-defined wavelengths, such as elevation, we assign default wavelengths equivalent to those of the RGB channels. First, a 1D sine-cosine positional encoding is applied to embed these wavelengths into a higher-dimensional space: $\mathbf{V}_{\lambda}=\text{PE}(\mathbf{\lambda})\in\mathbb{R}^{C\times d_{\lambda}},$ where $d_{\lambda}$ is the embedding dimension. The encoded wavelengths $\mathbf{V}_{\lambda}$ are then transformed using two fully-connected layers with residual connections. Next, a lightweight Transformer (Vaswani et al. 2017) encoder processes the transformed wavelength embeddings. Specifically, we concatenate the transformed embeddings $\mathbf{V^{\prime}}_{\lambda}$ with the learnable weight query tokens $\mathbf{Q}_{w}$ and a learnable bias query token $\mathbf{Q}_{b}$ as input to the Transformer encoder. After the Transformer layer, the output features are passed through two fully-connected layers to generate the dynamic convolutional weights $\mathbf{M}_{w}$ and biases $\mathbf{M}_{b}$. Finally, the dynamically generated weights and biases are used for patch embedding to handle EO data modalities with varying spectral channels.
Vision-models enhanced CLIP
To address RQ2, we propose VECT to enhance the spatial coherence of CLIP features by knowledge distillation from open vision models. Specifically, in addition to the contrastive loss, VECT leverages a feature-matching loss to align the intermediate features of DOFA-CLIP with features from multiple pre-trained vision models: SigLIP (Zhai et al. 2023), DINOv2 (Oquab et al. 2023), and ViT (Dosovitskiy et al. 2020). As illustrated in Fig. 3, we extract three RGB channels from multispectral or hyperspectral images as input to the teacher model. The student features $\mathbf{F}_{s}\in\mathbb{R}^{d_{t}\times H^{\prime}\times W^{\prime}}$ are then matched to the teacher’s features $\mathbf{F}_{t}\in\mathbb{R}^{d_{t}\times H^{\prime}\times W^{\prime}}$ using a combination of $L_{1}$ loss, cosine embedding loss, and mean squared error (MSE) loss:
DOFA-CLIP is trained on GeoLangBind-2M using both the pairwise sigmoid loss $L_{\text{siglip}}$ for contrastive learning and distillation loss for feature alignment. The total loss function is:
Modality-aware knowledge agglomeration
@@
Considering RQ3, we introduce the modality-aware knowledge agglomeration (MaKA) module, which is used to translate DOFA-CLIP features in a modality-aware manner to facilitate effective feature matching in VECT. MaKA comprises three key components: a wavelength-aware prompt generator, a modality-aware conditional layer normalization, and a feature projector.
1. Wavelength-aware prompt generator